The base of an isosceles triangle is 3 more than 2 times the length of the legs and the perimeter is 23. What is the length of the base?

Answer by lwsshak3(11628)

The base of an isosceles triangle is 3 more than 2 times the length of the legs and the perimeter is 23. What is the length of the base?
***
let x=length of legs
2x+3=length of base (3 more than 2 times the length of the legs)
2(length+width)=perimeter
2x+2x+3)=23
4x=20
X=5
2x+3=13
length of base=13 units